OpenStudy (anonymous):

Sam’s plasma TV has a 14 in by 18 in screen with a plastic trim border. The total area of the TV and border is 480 in2. How wide is the trim around the TV?

OpenStudy (anonymous):

tv area is 14 x 18 = 252. The total area - tv area is the trim area 480 - 252 = 228 480 = (14 + x)(18 + x) = x^2 + 32x + 252 x^2 + 32x - 228 = 0 You can now just solve for "x" using the quadratic formula:\[x = \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\]

OpenStudy (anonymous):

So, x = 6

OpenStudy (anonymous):

thank you very much. im pretty much a pro at the quadratic formula so i should get that answer in like 5 minutes

OpenStudy (anonymous):

That's good to hear. I almost wish that I didn't completely work out the final answer. Just cover it up what I did and use it for verification! :-)

OpenStudy (anonymous):

Good luck in all of your studies and thx for the recognition! And you're welcome! @alassen

OpenStudy (anonymous):

Ill reply back in a little bit to see if i got 6

OpenStudy (anonymous):

ok, I'll hang around.

OpenStudy (anonymous):

so far i have x = -32 + or - square root of 1936 divided by 2

OpenStudy (anonymous):

is that right so far?

OpenStudy (anonymous):

Yes, that's right, but you can eliminate the "-" in the +- because you will have a physical measurement which will be positive.

OpenStudy (anonymous):

now i have x = -32 + or - square root of 44 times square root of 44 divided by 2

OpenStudy (anonymous):

Yes, but again, you should eliminate the "- square root of..." It's going to be + the square root of... So, (44 - 32)/2

OpenStudy (anonymous):

i got 6! thanks a lot

OpenStudy (anonymous):

You're welcome! Cool that you got it!